# A new framework for nonlinear Kalman filters

### v6

@@ 2.3 The proposed new framework for nonlinear Kalman filters
1:Input: Process noise covariance matrix $\boldsymbol{Q}_{k}$, Measurement noise covariance matrix $\boldsymbol{R}_{k}$, state transition function $\boldsymbol{f}(\boldsymbol{x},\boldsymbol{u})$, measurement function $\boldsymbol{h}{(\boldsymbol{x})}$, system inputs $\boldsymbol{u}_{k}$, measurements $\boldsymbol{z}_{k}$
@@
3:$\boldsymbol{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\boldsymbol{x}_{0}]$
4:$\boldsymbol{P}_{0|0}=\operatorname{\mathbb{E}}[(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})^{T}]$
@@
7:     $\boldsymbol{F}_{k}=\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{x}}|_{\boldsymbol{\hat{x}}_{k-1|k-1},\boldsymbol{u}_{k-1}}$
8:     $\boldsymbol{\hat{x}}_{k|k-1}=\boldsymbol{f}(\boldsymbol{\hat{x}}_{k-1|k-1},\boldsymbol{u}_{k-1})$
9:     $\boldsymbol{P}_{k|k-1}=\boldsymbol{F}_{k}\boldsymbol{P}_{k-1|k-1}\boldsymbol{F}_{k}^{T}+\boldsymbol{Q}_{k}$
@@
13:         $\boldsymbol{H}_{k|k-1}=\frac{\partial\boldsymbol{h}}{\partial\boldsymbol{x}}|_{\boldsymbol{x}_{i-1}}$
14:         $\boldsymbol{\hat{y}}_{k|k-1}=\boldsymbol{h}(\boldsymbol{x}_{i-1})+\boldsymbol{H}_{k|k-1}(\boldsymbol{x}_{i-1}-\boldsymbol{\hat{x}}_{k|k-1})$
15:         $\boldsymbol{P}_{xy,k|k-1}=\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{k|k-1}^{T}$
16:         $\boldsymbol{P}_{y,k|k-1}=\boldsymbol{H}_{k|k-1}\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{k|k-1}^{T}$
17:         $\boldsymbol{S}_{k|k-1,i}=\boldsymbol{P}_{y,k|k-1}+\boldsymbol{R}_{k}$
18:         $\boldsymbol{K}_{i}=\boldsymbol{P}_{xy,k|k-1}\boldsymbol{S}_{k|k-1,i}^{-1}$
19:         $\boldsymbol{x}_{i}=\boldsymbol{\hat{x}}_{k|k-1}+\boldsymbol{K}_{i}(\boldsymbol{\boldsymbol{z}_{k}-\hat{y}}_{k|k-1})$
20:         if $i>1$ and $||\boldsymbol{x}_{i}-\boldsymbol{x}_{i-1}||>||\boldsymbol{x}_{i-1}-\boldsymbol{x}_{i-2}||$ then
21:              $\boldsymbol{S}_{k|k-1,i}=\boldsymbol{S}_{k|k-1,i-1}$
22:              $\boldsymbol{K}_{i}=\boldsymbol{K}_{i-1}$
23:              $\boldsymbol{x}_{i}=\boldsymbol{x}_{i-1}$
@@
30:     $\boldsymbol{S}_{k|k-1}=\boldsymbol{S}_{k|k-1,i}$
31:     $\boldsymbol{K}=\boldsymbol{K}_{i}$
32:     $\boldsymbol{\hat{x}}_{k|k}=\boldsymbol{x}_{i}$
33:     $\boldsymbol{P}_{k|k}=\boldsymbol{P}_{k|k-1}-\boldsymbol{K}\boldsymbol{S}_{k|k-1}\boldsymbol{K}^{T}$
@@
Since IKF also uses (9) to update the state covariance matrix, it can also be considered as another framework for nonlinear KFs that is similar to the conventional framework. A natural question is: can we combine the new framework with the IKF to get a higher estimation accuracy? Unfortunately, the answer is no because a prerequisite for the new framework to give accurate covariance estimation is that the approximations made at the predicted states ($P_{y,k|k-1},S_{k|k-1}$, and $P_{xy,k|k-1}$) are independent of the approximations made at the updated states ($P_{y,k|k},S_{k|k}$, and $P_{xy,k|k}$). However, after IKF converges, these two sets of approximations become the same, which violates this prerequisite in Theorem 2.5. For example, the IEKF algorithm shown in Algorithm 2 approximates the measurement function iteratively at $x_{i-1}$ (rows 9–13). When the algorithm converges, in the final iteration, we have $x_{i-1}\approx x_{i}=\hat{x}_{k|k}$, meaning that the approximation of the measurement functions in the final iteration will stay the same after the state update. In fact, to the best of our knowledge, IKF-related algorithms are the only existing types of nonlinear KFs that can not get large benefits from the proposed framework. Nevertheless, for any type of nonlinear KF that can be combined with IKF, its accuracy can be greatly improved when it is combined with our new framework instead. In this sense, we can still say that all existing types of nonlinear KFs can benefit from the proposed framework.
@@ 3 Results
To demonstrate the effectiveness of the proposed framework, we validated it in five different applications suitable for nonlinear KFs. A brief description of the five applications is presented in Table 1, and a detailed description of the system models and parameter setup can be found in Appendix 5. For each scenario, we combined the old and proposed framework with four different types of nonlinear KFs: EKF, EKF2, UKF, and CKF. We also implemented the IEKF for comparison. The detailed EKF, EKF2, UKF, and CKF algorithms used in the paper can be found in Algorithms 36 in Appendix 6. All these nine $(4+4+1)$ nonlinear KFs are used for state estimation under different measurement noise setups (the covariance of the measurement noises differs in different setups). To guarantee the fairness of the comparison and reduce the effect of randomness, each KF is simulated 10,000 times for each measurement noise setup, and all the KFs use the same random seed during the simulation. This means that the random initial states’ errors, random process noises, and random measurement noises are the same for the $i^{th}$ simulation of EKF and the $i^{th}$ simulation of UKF. In other words, all these random noises were generated for the 10,000 runs a priori and recalled for simulating each estimation method.
@@
As we mentioned in the Introduction, the fundamental idea of the proposed framework is to make a more accurate covariance estimation by re-approximating the measurement functions after the state update. To validate if the covariance estimation really becomes more accurate in the new framework, we compared the actual and estimated RMSE of the x-axis position (the first state in the system) in the old and new framework in Fig. 8. In the figure, for each type of nonlinear KFs (represented by different line colors) and framework (the old and new frameworks are respectively shown in the top and bottom subfigures), the closeness between the actual RMSE (the solid line) and estimated RMSE (dotted line) represents the accuracy of the state covariance estimation. Apparently, when the standard deviations of the two measurements are lower than 0.1 meters, the conventional framework always significantly underestimates the covariance, sometimes by several orders of magnitude, regardless of the type of nonlinear KFs used. On the other hand, for the new framework, only the RMSE of EKF is underestimated, and the covariance estimation of other types of nonlinear KFs matches the actual value very well. This, again, can be explained by the fact that EKF neglects the second and higher-order information of the measurement functions, so its covariance estimation is less accurate than other methods. Nevertheless, the covariance estimation of EKF is still much more precise when combined with the new framework.
@@
Another interesting observation of Fig. 8 is that the estimated RMSE of the x-axis position (the dashed lines) is very similar for different combinations of frameworks and KFs. This suggests that the accuracy of the estimation mainly relates to whether the covariance estimation is accurate or not, which leads to the conclusion mentioned at the end of the Introduction – to achieve a more accurate estimation, finding the optimal Kalman gain is not as important as accurately estimating the covariance.
@@
From the result, we can see that the new framework generally increases the computational time by 10–90%, varying for different nonlinear systems. Namely, when the system has linear state transition functions and nonlinear measurement functions (which is true for the first two applications in the table), the “update” step should be accountable for most of the computational time in the conventional framework and re-approximating the measurement functions can almost double the runtime in the worst case. The normalized runtime of the new framework is related to the type of nonlinear KFs as well. In UKF and CKF, one of the most time-consuming steps is to calculate the square root of the state covariance matrix. This step is respectively done one, two, two, and three times in each iteration of UKF (old framework), UKF (new framework), CKF (old framework), and CKF (new framework). Therefore, the computational time increase (in percentage) brought by the new framework is longer for UKF than for CKF. Noticing that the new framework can reduce the estimation error by more than an order of magnitude when the measurement noise is small, we can say that the extra computational complexity is generally worth it. Another thing to note is that the combination of EKF and the new framework requires significantly less computational time than the combination of any other nonlinear KFs and the old framework. At the same time, the former’s accuracy is generally higher than the latter in the applications we examined in this paper. Therefore, even in situations where the computational time is of major concern, the proposed framework can still be helpful because the new EKF can outperform the old EKF2, UKF, and CKF, in addition to saving a lot of computational time.
@@ 4 Conclusion
{appendices}
5 The detailed simulation setup
@@
The states are the target’s positions $(x_{1},x_{2},x_{3})$ and speeds $(v_{1},v_{2},v_{3})$ along the three axes. The measurements $y_{1},y_{2}$ are the distances between the target and the two sensors. The inputs are the target’s accelerations along the three axes, which are assumed to be zero in this example. When the time interval is set to $\Delta t=1$s, the system’s state transition and measurement functions can be written as (22,23). Note that the noises are omitted in the equations.
@@
The true initial states of the target are $x_{1,0}=10,x_{2,0}=-10,x_{3,0}=50,v_{1,0}=1,v_{2,0}=2,v_{3,0}=0$; the initial state covariance matrix $P_{0}$ is diagonal, and its diagonal elements are 100, 100, 100, 0.01, 0.01, and 0.01. The initial state estimations are generated randomly from the multivariate Gaussian distribution with covariance $P_{0}$ and mean equal to the true initial states. The process noise covariance matrix is also diagonal, and its diagonal values are 0, 0, 0, $10^{–6}$, $10^{–6}$, and $10^{–6}$.
5.2 Terrain-referenced navigation
This example is adapted from Example 3.3.2 in [22]. As shown in Fig. 11(b), a plane flies above an uneven landscape. The states are the planes’ positions along the two axes, denoted as $x_{1}$ and $x_{2}$. The measurement $y_{k}$ is the terrain’s elevation at the plane’s present location, which, in practice, is calculated as the difference between the barometer and radar altimeter measurements. The plane knows the contour map in advance, so the terrain’s elevation can tell the aircraft its possible locations. However, this task is difficult because all the points on a contour line have the same elevation, and system knowledge must be combined for accurate state estimations. The system inputs are the plane’s speeds along the x and y axes, which are 0.5 km/s and 0, respectively. When the time interval is set to $\Delta t=1$s, the system’s state transition and measurement functions can be written as (24, 25). Note that all the units for distance in this example are kilometers, and the noises are omitted in the equations.

### v10

@@ 2.4 Relationship to Iterated Kalman Filter
1:Input: Process noise covariance matrix $\bm{Q}_{k}$, Measurement noise covariance matrix $\bm{R}_{k}$, state transition function $\bm{f}(\bm{x},\bm{u})$, measurement function $\bm{h}{(\bm{x})}$, system inputs $\bm{u}_{k}$, measurements $\bm{z}_{k}$
@@
3:$\bm{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\bm{x}_{0}]$
4:$\bm{P}_{0|0}=\operatorname{\mathbb{E}}[(\bm{\hat{x}}_{0|0}-\bm{x}_{0})(\bm{\hat{x}}_{0|0}-\bm{x}_{0})^{T}]$
@@
7:     $\bm{F}_{k}=\frac{\partial\bm{f}}{\partial\bm{x}}|_{\bm{\hat{x}}_{k-1|k-1},\bm{u}_{k-1}}$
8:     $\bm{\hat{x}}_{k|k-1}=\bm{f}(\bm{\hat{x}}_{k-1|k-1},\bm{u}_{k-1})$
9:     $\bm{P}_{k|k-1}=\bm{F}_{k}\bm{P}_{k-1|k-1}\bm{F}_{k}^{T}+\bm{Q}_{k}$
@@
13:         $\bm{H}_{k|k-1}=\frac{\partial\bm{h}}{\partial\bm{x}}|_{\bm{x}_{i-1}}$
14:         $\bm{\hat{y}}_{k|k-1}=\bm{h}(\bm{x}_{i-1})+\bm{H}_{k|k-1}(\bm{x}_{i-1}-\bm{\hat{x}}_{k|k-1})$
15:         $\bm{P}_{xy,k|k-1}=\bm{P}_{k|k-1}\bm{H}_{k|k-1}^{T}$
16:         $\bm{P}_{y,k|k-1}=\bm{H}_{k|k-1}\bm{P}_{k|k-1}\bm{H}_{k|k-1}^{T}$
17:         $\bm{S}_{k|k-1,i}=\bm{P}_{y,k|k-1}+\bm{R}_{k}$
18:         $\bm{K}_{i}=\bm{P}_{xy,k|k-1}\bm{S}_{k|k-1,i}^{-1}$
19:         $\bm{x}_{i}=\bm{\hat{x}}_{k|k-1}+\bm{K}_{i}(\bm{z}_{k}-\bm{\hat{y}}_{k|k-1})$
20:         if $i>1$ and $||\bm{x}_{i}-\bm{x}_{i-1}||>||\bm{x}_{i-1}-\bm{x}_{i-2}||$ then
21:             $\bm{S}_{k|k-1,i}=\bm{S}_{k|k-1,i-1}$
22:             $\bm{K}_{i}=\bm{K}_{i-1}$
23:             $\bm{x}_{i}=\bm{x}_{i-1}$
@@
30:     $\bm{S}_{k|k-1}=\bm{S}_{k|k-1,i}$
31:     $\bm{K}=\bm{K}_{i}$
32:     $\bm{\hat{x}}_{k|k}=\bm{x}_{i}$
33:     $\bm{P}_{k|k}=\bm{P}_{k|k-1}-\bm{K}\bm{S}_{k|k-1}\bm{K}^{T}$
@@
Since IKF also uses (10) to update the state covariance matrix, it can also be considered as another framework for nonlinear KFs that is similar to the conventional framework. A natural question is: can we combine the new framework with the IKF to get a higher estimation accuracy? Unfortunately, the answer is no because a prerequisite for the new framework to give accurate covariance estimation is that the approximations made at the predicted states (${\color[rgb]{1,0,0}P_{y,k|k-1}},{\color[rgb]{1,0,0}S_{k|k-1}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$) are independent of the approximations made at the updated states (${\color[rgb]{1,0,0}P_{y,k|k}},{\color[rgb]{1,0,0}S_{k|k}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k}}$). However, after IKF converges, these two sets of approximations become the same, which violates this prerequisite in Theorem 2. For example, the IEKF algorithm shown in Algorithm 2 approximates the measurement function iteratively at $x_{i-1}$ (rows 9–13). When the algorithm converges, in the final iteration, we have $x_{i-1}\approx x_{i}=\hat{x}_{k|k}$, meaning that the approximation of the measurement functions in the final iteration will stay the same after the state update. In fact, to the best of our knowledge, IKF-related algorithms are the only existing types of nonlinear KFs that can not get large benefits from the proposed framework. Nevertheless, for any type of nonlinear KF that can be combined with IKF, its accuracy can be greatly improved when it is combined with our new framework instead. In this sense, we can still say that all existing types of nonlinear KFs can benefit from the proposed framework.
@@ 3 Results
To demonstrate the effectiveness of the proposed framework, we validated it in five different applications suitable for nonlinear KFs. A brief description of the five applications is presented in Table 1, and a detailed description of the system models and parameter setup can be found in Appendix B. For each scenario, we combined the old and proposed framework with four different types of nonlinear KFs: EKF, EKF2, UKF, and CKF. We also implemented the IEKF for comparison. The detailed EKF, EKF2, UKF, and CKF algorithms used in the paper can be found in Algorithms 35 in Appendix C. All these nine $(4+4+1)$ nonlinear KFs are used for state estimation under different measurement noise setups (the covariance of the measurement noises differs in different setups). To guarantee the fairness of the comparison and reduce the effect of randomness, each KF is simulated 10,000 times for each measurement noise setup, and all the KFs use the same random seed during the simulation. This means that the random initial states’ errors, random process noises, and random measurement noises are the same for the $i^{th}$ simulation of EKF and the $i^{th}$ simulation of UKF. In other words, all these random noises were generated for the 10,000 runs a priori and recalled for simulating each estimation method.
@@
As we mentioned in the Introduction, the fundamental idea of the proposed framework is to make a more accurate covariance estimation by re-approximating the measurement functions after the state update. To validate if the covariance estimation really becomes more accurate in the new framework, we compared the actual and estimated RMSE of the x-axis position (the first state in the first system) in the old and new framework in Fig. 8. In the figure, for each type of nonlinear KF (represented by different line colors) and framework (the old and new frameworks are respectively shown in the top and bottom subfigures), the closeness between the actual RMSE (the solid line) and estimated RMSE (dotted line) represents the accuracy of the state covariance estimation. Apparently, when the standard deviations of the two measurements are lower than 0.1 meters, the conventional framework always significantly underestimates the covariance, sometimes by several orders of magnitude, regardless of the type of nonlinear KFs used. On the other hand, for the new framework, only the RMSE of EKF is underestimated, and the covariance estimation of other types of nonlinear KFs matches the actual value very well. This, again, can be explained by the fact that EKF neglects the second and higher-order information of the measurement functions, so its covariance estimation is less accurate than other methods. Nevertheless, the covariance estimation of EKF is still much more precise when combined with the new framework.
@@
Another interesting observation of Fig. 8 is that the estimated RMSE of the x-axis position (the dashed lines) is very similar for different combinations of frameworks and KFs. This suggests that the accuracy of the estimation mainly relates to whether the covariance estimation is accurate or not, which leads to the conclusion mentioned at the end of the Introduction — to achieve a more accurate estimation, finding the optimal Kalman gain is not as important as accurately estimating the covariance.
@@
From the result, we can see that the new framework generally increases the computational time by 10–90%, varying for different nonlinear systems. Namely, when the system has linear state transition functions and nonlinear measurement functions (which is true for the first two applications in the table), the “update” step should be accountable for most of the computational time in the conventional framework and re-approximating the measurement functions can almost double the runtime in the worst case. However, noticing that the new framework can reduce the estimation error by more than an order of magnitude when the measurement noise is small, we can say that the extra computational complexity is generally worth it. Another thing to note is that the combination of EKF and the new framework requires significantly less computational time than the combination of any other nonlinear KFs and the old framework. At the same time, the former’s accuracy is generally higher than the latter in the applications we examined in this paper. Therefore, even in situations where the computational time is of major concern, the proposed framework can still be helpful because the new EKF can outperform the old EKF2, UKF, and CKF, in addition to saving a lot of computational time.
@@ 4 Conclusion
{ack}
Appendix A Supplementary proof for Theorem 1
Lemma 3.
Let $(\Omega,\mathcal{F},\mathbb{P})$ be an underlying probability space. Assume the random column vectors
$x:\Omega\to\mathbb{R}^{n},\;y:\Omega\to\mathbb{R}^{n}$
satisfy $\mathbb{E}[||x||^{2}]<\infty$ and $\mathbb{E}[||y||^{2}]<\infty$ (i.e., $x,y\in L^{2}(\Omega;\mathbb{R}^{n})$). Then, the operation on $L^{2}(\Omega;\mathbb{R}^{n})$ by
$f(u,v)=\mathbb{E}[u^{T}v]$
is an inner product on $L^{2}(\Omega;\mathbb{R}^{n})$.
{pf}
We start by proving that $f$ is well-defined on $L^{2}$. For any $u,v\in L^{2}$, both $\mathbb{E}[u^{T}u]$ and $\mathbb{E}[v^{T}v]$ are finite. Therefore, $\mathbb{E}[(u+v)^{T}(u+v)]$ should also be finite. Notice that
$\mathbb{E}[(u+v)^{T}(u+v)]=\mathbb{E}[u^{T}u]+\mathbb{E}[v^{T}v]+2\mathbb{E}[u^{T}v],$
we can conclude that $\mathbb{E}[u^{T}v]$ should also be finite. Therefore, $f$ is well-defined on $L^{2}$. Next, to prove that $f$ is an inner product, we need to show that it satisfies symmetry, linearity in the first argument, and positive-definiteness. Symmetry is satisfied because $\mathbb{E}[u^{T}v]=\mathbb{E}[v^{T}u]$. Linearity is satisfied because for any scalars $a,b$ and $x,y,z\in L^{2}$,
$\mathbb{E}[(ax+by)^{T}z]=a\mathbb{E}[x^{T}z]+b\mathbb{E}[y^{T}z].$
Positive-definiteness is satisfied because for any $x\in L^{2}$, $\mathbb{E}[x^{T}x]\geq 0$, and the equality sign only holds when $x\equiv 0$.
Lemma 4.
Let ${\color[rgb]{1,0,0}S}$ be a random symmetric and positive definite matrix whose mean is the identity matrix $I$. Let $\bm{v}$ be a random vector whose mean is $\bm{v}_{0}$. The following inequality holds:
$\mathbb{E}(\bm{v}^{T}{\color[rgb]{1,0,0}S^{-1}}\bm{v})\geq\bm{v}_{0}^{T}\bm{v}_{0}$ (22)
Note that the equality sign only holds when ${\color[rgb]{1,0,0}S^{-1}}\bm{v}$ is constant.
{pf}
Let $\bm{x}={\color[rgb]{1,0,0}S^{-\frac{1}{2}}}\bm{v},\bm{y}={\color[rgb]{1,0,0}S^{\frac{1}{2}}}\bm{v}_{0}$. From the definitions,
$\begin{cases}\mathbb{E}[\bm{x}^{T}\bm{x}]=\mathbb{E}[\bm{v}^{T}{\color[rgb]{1,0,0}S^{-1}}\bm{v}]\\ \mathbb{E}[\bm{y}^{T}\bm{y}]=\mathbb{E}[\bm{v}_{0}^{T}{\color[rgb]{1,0,0}S}\bm{v}_{0}]=\bm{v}_{0}^{T}\bm{v}_{0}\\ \mathbb{E}[\bm{x}^{T}\bm{y}]=\mathbb{E}[\bm{v}^{T}\bm{v}_{0}]=\bm{v}_{0}^{T}\bm{v}_{0}\end{cases}$ (23)
According to Lemma 3, $\mathbb{E}[\bm{x}^{T}\bm{y}]$ is an inner product on $L^{2}(\Omega;\mathbb{R}^{n})$. Applying Cauchy–Schwarz inequality, we have:
$\mathbb{E}[\bm{v}^{T}{\color[rgb]{1,0,0}S^{-1}}\bm{v}]\geq\frac{(\bm{v}_{0}^{T}\bm{v}_{0})^{2}}{\bm{v}_{0}^{T}\bm{v}_{0}}$ (24)
which completes the proof. Note that the equality holds only when ${\color[rgb]{1,0,0}S^{-1}}\bm{v}$ is constant.
Appendix B The detailed simulation setup
@@
The states are the target’s positions $(x_{1},x_{2},x_{3})$ and speeds $(v_{1},v_{2},v_{3})$ along the three axes. The measurements $y_{1},y_{2}$ are the distances between the target and the two sensors. The inputs are the target’s accelerations along the three axes, which are assumed to be zero in this example. When the time interval is set to $\Delta t=1$s, the system’s state transition and measurement functions can be written as (26,27). Note that the noises are omitted in the equations.
@@
The true initial states of the target are $x_{1,0}=10,x_{2,0}=-10,x_{3,0}=50,v_{1,0}=1,v_{2,0}=2,v_{3,0}=0$; the initial state covariance matrix $P_{0}$ is diagonal, and its diagonal elements are 100, 100, 100, 0.01, 0.01, and 0.01. The initial state estimations are generated randomly from the multivariate Gaussian distribution with covariance $P_{0}$ and mean equal to the true initial states. The process noise covariance matrix is also diagonal, and its diagonal values are 0, 0, 0, $10^{-6}$, $10^{-6}$, and $10^{-6}$.
B.2 Terrain-referenced navigation
This example is adapted from Example 3.3.2 in [26]. As shown in Fig. 11(b), a plane flies above an uneven landscape. The states are the planes’ positions along the two axes, denoted as $x_{1}$ and $x_{2}$. The measurement $y_{k}$ is the terrain’s elevation at the plane’s present location, which, in practice, is calculated as the difference between the barometer and radar altimeter measurements. The plane knows the contour map in advance, so the terrain’s elevation can tell the aircraft its possible locations. However, this task is difficult because all the points on a contour line have the same elevation, and system knowledge must be combined for accurate state estimations. The system inputs are the plane’s speeds along the x and y axes, which are 0.5 km/s and 0, respectively. When the time interval is set to $\Delta t=1$s, the system’s state transition and measurement functions can be written as (28, 29). Note that all the units for distance in this example are kilometers, and the noises are omitted in the equations.